# BitMoD: Bit-serial Mixture-of-Datatype LLM Acceleration

### v1

@@ II-C Motivation
In order to further reduce the quantization error, recent LLM quantization studies adopt the per-group granularity [15, 19, 29, 18]. The per-group quantization further divides a weight channel $W_{f}^{1\times D}$ into $D/G$ groups, each with a group size of $G$. The group size introduces extra overhead to store the quantization parameters, i.e., scaling factor (and zero-point) for every group, and is usually set to 128 in SOTA quantization frameworks to balance accuracy and memory overhead [19, 29]. Fig. 2 demonstrates the benefits of per-group quantization by showing the maximum value and range in four representative LLMs at different granularity. The per-group granularity has the lowest maximum value and range, hence will have a lower quantization error compared to the other two granularity. Therefore, we focus on per-group quantization in this work.
@@ III-A Asymmetric FP3 and FP4 Data Types
TABLE III: Comparison between BitMoD and SOTA co-design frameworks for LLM acceleration
@@ IV BitMoD Hardware Accelerator
In this section, we describe the BitMoD hardware design, which leverages the bit-serial computing paradigm to offer a good trade-off between weight precision, model accuracy and hardware efficiency. Section IV-A develops a unified bit-serial representation of different low-precision data types supported by BitMoD. Section IV-B details the microarchitecture of the BitMoD processing element (PE). Section IV-C presents the overall accelerator architecture.
@@ IV-B BitMoD Processing Element
While BitMoD is able to quantize weight to low precision, activation still remains in FP16. To address this challenge, we propose a mixed-precision bit-serial PE as shown in Fig. 5. In every cycle, the PE performs a 4-way dot product between four bit-serial weight terms ($w$) and four FP16 activations ($a$). Step  1 first aligns the sum of exponents ($a_{e}+w_{e}$) to compute the delta exponent ($\delta_{e}$). It also generates the sign ($y_{s}$) of every product between a weight term and activation. Step  2 performs the bit-serial multiplication between the 1-bit weight mantissa ($w_{m}$) and 11-bit activation mantissa ($a_{m}$) including the hidden bit. The multiplication result is aligned by a right-shifter that is controlled by the delta exponent. We reserve 3 extra bits in the shifter result to account for rounding to the nearest even as suggested by Awad et al. [4]. The bit-serial dot product of the mantissa is then computed using an adder tree. After producing the bit-serial dot product, Step  3 performs accumulation by first multiplying the dot product with the weight bit-significance ($W_{bsig}$), followed by adding with the accumulator mantissa ($m_{ACC}$). The accumulated mantissa is then normalized to update the accumulator exponent ($e_{ACC}$). Since BitMoD adopts per-group quantization, the accumulated group partial sum must be dequantized on the fly. To reduce this hardware cost, Step  4 performs dequantization in a bit-serial manner. Specifically, the accumulator mantissa is multiplied by one bit of the group scaling factor ($\Delta_{i}$) in every cycle, followed by shift-and-add to obtain the exponent ($e_{GRP}$) and mantissa ($e_{GRP}$) of the dequantized partial sum.
@@ V-A Experimental Methodology
TABLE VII: Accuracy (higher is better) of discriminative tasks using different data types under per-group weight quantization.
@@
Figure 7: Speedup of different accelerators (higher is better).
Figure 8: Energy consumption breakdown of different accelerators ($\downarrow$). “LL” and “LY” stand for ‘lossless” and ‘lossy”, respectively.
@@ Acknowledgment
This project is supported in part by Intel Corporation, the National Science Foundation under Grant No. 2339084, and the Engineering and Physical Sciences Research Council (EPSRC) under Grant No. EP/S030069/1. We would like to thank Mahesh Iyer, Grace Zgheib, Sergey Gribok, and the anonymous reviewers for their constructive feedback.

### v2

@@ II-C Motivation
In order to further reduce the quantization error, recent LLM quantization studies adopt the per-group granularity [16, 20, 30, 19]. The per-group quantization further divides a weight channel $W_{f}^{1\times D}$ into $D/G$ groups, each with a group size of $G$. The group size introduces extra overhead to store the quantization parameters, i.e., scaling factor (and zero-point) for every group, and is usually set to 128 in SOTA quantization frameworks to balance accuracy and memory overhead [20, 30]. Fig. 2 demonstrates the benefits of per-group quantization by showing the maximum value and range in four representative LLMs at different granularities. The per-group granularity has the lowest maximum value and range, hence will have a lower quantization error compared to the other two granularity. Therefore, we focus on per-group quantization in this work.
@@ III-A Asymmetric FP3 and FP4 Data Types
TABLE III: Comparison between BitMoD and SOTA co-design frameworks for quantized LLM acceleration.
@@ IV BitMoD Hardware Accelerator
In this section, we describe the BitMoD hardware design, which leverages the bit-serial computing paradigm to offer a good trade-off between weight precision, model accuracy and hardware efficiency. Section IV-A develops a unified bit-serial representation of different low-precision data types supported by BitMoD. Section IV-B details the microarchitecture of the BitMoD processing element (PE). Finally, section IV-C presents the overall accelerator architecture.
@@ IV-B BitMoD Processing Element
While BitMoD is able to quantize weight to low precision, activation still remains in FP16. To address this challenge, we propose a mixed-precision bit-serial PE as shown in Fig. 5. In every cycle, the PE performs a 4-way dot product between four bit-serial weight terms ($w$) and four FP16 activations ($a$). Step  1 first aligns the sum of exponents ($a_{e}+w_{e}$) to compute the delta exponent ($\delta_{e}$). It also generates the sign ($y_{s}$) of every product between a weight term and activation. Step  2 performs the bit-serial multiplication between the 1-bit weight mantissa ($w_{m}$) and 11-bit activation mantissa ($a_{m}$) including the hidden bit. The multiplication result is aligned by a right-shifter that is controlled by the delta exponent. We reserve 3 extra bits in the shifter result to account for rounding to the nearest even as suggested by Awad et al. [5]. The bit-serial dot product of the mantissa is then computed using an adder tree. After producing the bit-serial dot product, Step  3 performs accumulation by first multiplying the dot product with the weight bit-significance ($W_{bsig}$), followed by adding with the accumulator mantissa ($m_{ACC}$). The accumulated mantissa is then normalized to update the accumulator exponent ($e_{ACC}$). Since BitMoD adopts per-group quantization, the accumulated group partial sum must be dequantized on the fly. To reduce this hardware cost, Step  4 performs dequantization in a bit-serial manner. Specifically, the accumulator mantissa is multiplied by one bit of the group scaling factor ($\Delta_{i}$) in every cycle, followed by shift-and-add to obtain the exponent ($e_{GRP}$) and mantissa ($m_{GRP}$) of the dequantized partial sum.
@@ V-A Experimental Methodology
TABLE VII: Accuracy ($\uparrow$) of discriminative tasks using different data types under per-group weight quantization.
@@
Figure 7: Speedup ($\uparrow$) of different accelerators.
Figure 8: Energy consumption ($\downarrow$) of different accelerators. “LL” and “LY” stand for ‘lossless” and ‘lossy”, respectively.
@@ Acknowledgment
This project is supported in part by Intel Corporation, the National Science Foundation under Grant No. 2339084, and the Engineering and Physical Sciences Research Council (EPSRC) under Grant No. EP/S030069/1. We would like to thank Jordan Dotzel, Hongxiang Fan, Stylianos I. Venieris, Alexandros Kouris, Mahesh Iyer, Grace Zgheib, Sergey Gribok, and the anonymous reviewers for their constructive feedback.
Appendix A Artifact Appendix
A-A Abstract
This artifact appendix describes how to access and evaluate the BitMoD artifact [2] to reproduce the experiments as performed in Section V.
A-B Artifact Check-List (Meta-Information)
  • Program: Python, Shell script
  • Runtime Environment: Ubuntu 20.04
  • Hardware: NVIDIA GPU with $\geq$40 GB of VRAM (e.g., A6000).
  • Model (LLM): OPT-1.3B, Phi-2B, Yi-6B, Llama-2-7B, Llama-2-13B, Llama-3-8B.
  • Dataset: Wikitext, C4.
  • Experiments: LLM perplexity evaluation in Section V-B and Section V-E. Accelerator performance evaluation in Section V-C.
  • Disk space required (approximately): 512 GB.
  • How much time is needed to complete experiments (approximately)?: 50 hours.
  • Publicly available?: Yes.
  • Archived: https://doi.org/10.5281/zenodo.14252531
A-C Description
A-C1 How to Access
We maintain a publicly-available repository [2] where we have open-sourced all of our artifacts.
A-C2 Hardware Dependencies
One NVIDIA GPU with at least 40 GB of VRAM (e.g., A6000), in addition to a normal desktop computer with at least 512 GB of free disk space.
A-C3 Software Dependencies
All experiments require Conda for managing virtual Python environments. The quantization experiments require CUDA. Other requirements are automatically installed by scripts in the following sections. When running experiments, please use a tmux session to make sure long-running jobs are not killed unexpectedly.
A-D Installation
For artifact evaluation, begin by downloading the top-level repository from Zenodo:
\$ wget -O BitMoD.zip https://zenodo.org/records/14252531/files/BitMoD-HPCA-25.zip
\$ unzip BitMoD.zip
The artifact is inside the unzipped BitMoD-HPCA-25 repository, which contains five sub-folders, each targeting one set of experiments:
  1. 1.
    bitmod-quant, which runs the baseline weight-only quantization with different data types. This can reproduce the results in Table VI and Table VIII.
  2. 2.
    bitmod-sim, contains a custom simulator that calculates the latency and energy of the BitMoD accelerator. This can reproduce the results in Fig. 7 and Fig. 8.
  3. 3.
    AWQ-BitMoD, which runs AWQ [30] with integer and BitMoD data types. This can reproduce the AWQ results in Table XI.
  4. 4.
    OmniQuant-BitMoD, which runs OmniQuant [42] with integer and BitMoD data types. This can reproduce the OmniQuant results in Table XI.
  5. 5.
    SmoothQuant-BitMoD, which runs SmoothQuant [52] with integer and BitMoD data types for weight quantization. This can reproduce the results in Table XII.
Please go to every sub-folder and refer to the corresponding ‘README.md’ for detailed setup instructions. Note that AWQ, OmniQuant, and SmoothQuant will require different conda environments. Hence, please change back to the base environment after installing a conda environment before creating the next.
For example, the AWQ environment can be created by running:
\$ cd AWQ-BitMoD
\$ conda create -n awq-bitmod python=3.10 -y
\$ conda activate awq-bitmod
# Follow ‘README.md’ inside AWQ-BitMoD folder to set up other dependencies.
\$ conda deactivate # change back to the base environment
The OmniQuant and SmoothQuant environments can be created in a similar way by repeating the above step and following their ‘README.md’ to set up other dependencies.
A-E Experiment workflow
Once the environment is set up, we will conduct five sets of experiments, each corresponding to one of the five sub-folders within the BitMoD-HPCA-25 repository.
A-E1 BitMoD Weight-only Quantization
Run the basic LLM weight-only quantization experiments to reproduce the results in Table VI and Table VIII.
\$ cd bitmod_quant
\$ conda activate awq-bitmod
In ‘run_exp.sh’, modify the ‘export’ command by specifying the HuggingFace home directory, ‘HF_HOME’, on your computer. By default, this can be set to your home directory.
\$ export HF_HOME=”your/HF_HOME/directory”
Then run the following:
\$ bash run_exp.sh
The perplexity result will be saved in the folder called ‘results_quant’.
A-E2 BitMoD Hardware Simulation
Before running the simulator, go to ‘bitmod_sim’ of the repository:
\$ cd bitmod_sim
\$ conda activate awq-bitmod
In ‘run_shape_profile.sh’, modify the ‘export’ command by specifying the HuggingFace home directory, ‘HF_HOME’, on your computer:
\$ export HF_HOME=”your/HF_HOME/directory”
Then generate the model shape information that can be passed to the accelerator simulator:
\$ bash run_shape_profile.sh
Next, run different simulators for the baseline FP16 accelerator, ANT, OliVe, and BitMoD:
\$ python test_baseline.py is_generation
\$ python test_ant.py is_generation
\$ python test_olive.py is_generation
\$ python test_bitmod.py is_generation is_lossless
The flag –is_generation is optional. When enabled / disabled, it will evaluate the hardware performance of generative / discriminative tasks. The flag –is_lossless is optional for BitMoD. When enabled / disabled, it will evaluate the hardware performance of lossless / lossy BitMoD quantization.
Finally, to generate Fig. 7 and Fig. 8 of the paper, go to ‘bitmod_sim/plot’ directory and run the Jupyter notebooks inside. Note that the cycle and energy numbers are the same as those output by the simulators.
A-E3 AWQ
Go to the ‘AWQ-BitMoD’ directory:
\$ cd AWQ-BitMoD
\$ conda activate awq-bitmod
In ‘run_awq.sh’ and ‘run_eval_ppl.sh’, modify the first ‘export’ command by specifying the HuggingFace home directory, ‘HF_HOME’, on your computer:
\$ export HF_HOME=”your/HF_HOME/directory”
Then, run the following two commands separately:
\$ bash run_awq.sh # will take several hours
\$ bash run_eval_ppl.sh
The perplexity results will be saved in the folder called ‘results’. You can compare these with the AWQ results in Table XI.
A-E4 OmniQuant
Go to the ‘OmniQuant-BitMoD’ directory:
\$ cd OmniQuant-BitMoD
\$ conda activate omniquant-bitmod
The comprehensive scripts to reproduce the Table XI OmniQuant results are available in the ‘scripts’ directory. Before running any command in the scripts, execute the following ‘export’ command and specify the HuggingFace home directory, ‘HF_HOME’, on your computer:
\$ export HF_HOME=”your/HF_HOME/directory”
In every shell script, you need to change the parameter of –model flag to the LLM path in your computer. By default, this can be set to the LLM directory from the official HuggingFace website. For example, inside the script ‘llama-2-13b-int.sh’, the Llama-2-7B model path can be specified with:
model meta-llama/Llama-2-7b-hf
After changing the –model flag to the correct model path, copy and execute every script’s python command under the ‘OmniQuant-BitMoD’ directory. You may check the perplexity results at the end of the log file specified by the –output_dir flag in every command, and compare those with the OmniQuant results in Table XI.
A-E5 SmoothQuant
Go to the ‘SmoothQuant-BitMoD’ directory:
\$ cd SmoothQuant-BitMoD
\$ conda activate smoothquant-bitmod
In ‘run_experiments.sh’, modify the ‘export’ command by specifying the HuggingFace home directory, ‘HF_HOME’, on your computer:
\$ export HF_HOME=”your/HF_HOME/directory”
Then, run the following command:
\$ bash run_experiments.sh
The perplexity results will be saved in the folder called ‘results_mod’. You can compare these results with the SmoothQuant results in Table XII.
A-F Evaluation and expected results
There are five result folders after running the above experiments:
  1. 1.
    bitmod-quant/results_quant, contains the perplexity results in Table VI and Table VIII.
  2. 2.
    bitmod-sim/plot, contains two Jupyter notebooks to reproduce Fig. 7 and Fig. 8, respectively.
  3. 3.
    AWQ-BitMoD/results, contains the AWQ results in Table XI.
  4. 4.
    OmniQuant-BitMoD/log, contains the OmniQuant results in Table XI.
  5. 5.
    SmoothQuant-BitMoD/results_mod, contains the SmoothQuant results in Table XII.
A-G Methodology
Submission, reviewing and badging methodology: